40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Mon Aug 15 10:27:56 2022, Koji, Update, General, Recap of the additional measures for the outage prep    
|
Mon Aug 15 15:43:49 2022, Anchal, Update, General, Complete power shutdown and startup documentation
|
Mon Aug 15 18:06:07 2022, Anchal, Update, General, c1vac issues, 1 pressure gauge died
|
Mon Aug 15 20:09:18 2022, Koji, Update, General, c1vac issues, 1 pressure gauge died
|
Wed Aug 17 10:23:05 2022, Tega, Update, General, c1vac issues, pressure gauge replacement
|
Fri Nov 18 09:00:27 2022, JC, Update, VAC, Pressure Gauge Information
|
|
Message ID: 17079
Entry time: Mon Aug 15 10:27:56 2022
Reply to this: 17080
|
Author: |
Koji |
Type: |
Update |
Category: |
General |
Subject: |
Recap of the additional measures for the outage prep |
|
|
[Yuta Koji]
(Report on Aug 12, 2022)
We went around the lab for the final check. Here are the additional notes.
- 1X9: The x-end frontend machine still had the AC power. The power strip to which the machine is connected was disconnected from the AC at the side of the rack. (Attachment 1)
- 1X8: The vacuum rack still supplied the AC to c1vac. This was turned off at the UPS. (Attachment 2)
- 1X6: VMI RFM hub still had the power. This was turned off at the rear switch. (Attachment 3)
- PSL: The PSL door was open (reported above). Closed. (Attachment 4)
- 1Y2: The LSC rack still had the DC power. The supplies were turned off at the KEPCO rack (the short rack). (Attachment 5)
Note that the top-right supply for the +15V is not used. (The one in the empty slot got busted). We may need some attention to the left-most one in the second row. It indicated a negative current. Is this just the current meter problem or is the supply broken?
- Control room: The CAD WS was turned off.
I declare that now we are ready for the power outage. |
|
|
|
|
|
|
|
|
|